angelcook2007:

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0

Florisalreadytaken:

a) the answer to that would be the discriminant, or \( b^{2}-4ac \) only real discriminant is the sign indicating that the quadratic is factorable. b) \( 4x^{2} + 6x + 2 = 0 \) -- find the value of the discriminant in this formula. we can see that: \[ a=4 \ \ \ \ \ ; \ \ \ \ \ b=6 \ \ \ \ \ c=2 \] putting that information into \( b^2-4ac \) we get: \[ D=6^{2}-4(4)(2) \] \[ D=36-16 \times 2 \] \[ D=36-32 \] \[ D=4 \]

angelcook2007:

@florisalreadytaken wrote:
a) the answer to that would be the discriminant, or \( b^{2}-4ac \) only real discriminant is the sign indicating that the quadratic is factorable. b) \( 4x^{2} + 6x + 2 = 0 \) -- find the value of the discriminant in this formula. we can see that: \[ a=4 \ \ \ \ \ ; \ \ \ \ \ b=6 \ \ \ \ \ c=2 \] putting that information into \( b^2-4ac \) we get: \[ D=6^{2}-4(4)(2) \] \[ D=36-16 \times 2 \] \[ D=36-32 \] \[ D=4 \]
I did what I understood and got 4 but I didn't know if it was right

Florisalreadytaken:

well, congrats on being correct then : )

angelcook2007:

lol

angelcook2007:

thank you, I had help on a similar question, but I forgot what I was supposed to do, I just went off of my faint memory on it